# A General Control-Theoretic Approach for Reinforcement Learning: Theory and Algorithms

### v1

@@ Appendix A Control-Based Reinforcement Learning Approach
Intuitively, Lemma A.1 shows that, for any policy families ${\mathbb{F}}_{1}\subset{\mathbb{F}}_{2}$ sufficiently close to each other, the fixed points $\tilde{q}_{1}$ and $\tilde{q}_{2}$ of the corresponding operators ${{\mathbf{T}}}_{1}$ and ${{\mathbf{T}}}_{2}$ are also close to each other. When the policy families ${\mathbb{F}}_{1}\subset{\mathbb{F}}_{2}$ are sufficiently rich and approach ${\mathbb{F}}$, then the fixed points of the corresponding operators ${{\mathbf{T}}}_{1}$ and ${{\mathbf{T}}}_{2}$ approach the unique fixed point of ${\mathbb{F}}$ satisfying (3), and therefore they approach the optimal $q$-value as promised by Bellman from Lemma 2.1. We formally characterize this asymptotic convergence of approximate optimality to global optimality in Theorem 2.1.
Theorem 2.1. Assume the state and action spaces are compact and ${\mathcal{F}}_{p}$ is uniformly continuous for each $p$. Consider a sequence of families of policy functions ${\mathbb{F}}_{1}\subset{\mathbb{F}}_{2}\subset\cdots\subset{\mathbb{F}}_{k-1}\subset{\mathbb{F}}_{k}$ such that $\bigcup_{i=1}^{k}{\mathbb{F}}_{i}\rightarrow{\mathbb{F}}$ as $k\rightarrow\infty$, with ${{\mathbb{S}}}$ and ${{\mathbb{S}}}_{i}$ respectively denoting the parameter sets corresponding to ${\mathbb{F}}$ and ${\mathbb{F}}_{i}$, $i\in[k]$. Then, $\sup_{x\in{{\mathbb{X}}}}\|\tilde{q}_{k}^{\prime}-\tilde{q}^{\prime}\|\rightarrow 0$ as $k\rightarrow\infty$.
@@
First, in the definition of the contraction mapping ${{\mathbf{T}}}$, the result of the supremum depends continuously on the set ${{\mathbb{S}}}$ and thus ${{\mathbf{T}}}$ depends continuously on ${{\mathbb{S}}}$. Next, since the state and action spaces are compact and ${\mathcal{F}}_{p}$ is uniformly continuous for each $p$, then [30, Lemma 2.2] shows that the unique fixed point depends continuously on the contraction mapping ${{\mathbf{T}}}$ and thus we find that $\tilde{q}^{\prime}$ depends continuously on the sets ${{\mathbb{S}}}_{i}$. This implies Lemma A.1, from which Theorem 2.1 then follows directly. ∎
A.3 Proof of Theorem 2.2.
Theorem 2.2. Consider a family of control policy functions ${\mathbb{F}}$ and its parameter set ${{\mathbb{S}}}$ with contraction operator ${{\mathbf{T}}}$ in the form of (2). We then have
$\nabla_{p}V_{{{\mathcal{F}}_{p}}}(x_{0})=\sum_{x\in{{\mathbb{X}}}}\>\sum_{k=0}^{\infty}\gamma^{k}\left(\mathbb{P}_{{\mathcal{F}}_{p}}(x_{0},x,k)\right)\>\sum_{u\in{\mathbb{U}}}\frac{\partial{\mathcal{F}}_{p,u}(x)}{\partial{\mathcal{F}}_{p}}\frac{\partial{\mathcal{F}}_{p}(x)}{\partial p}\tilde{q}_{u}(x,{{\mathcal{F}}}_{p}),$ (4)
where it is assumed that ${\mathcal{F}}_{p,u}(x)$ is differentiable w.r.t. ${\mathcal{F}}_{p}$ and ${\mathcal{F}}_{p}(x)$ is differentiable w.r.t. $p$, i.e., both $\partial{\mathcal{F}}_{p,u}(x)/\partial{\mathcal{F}}_{p}$ and $\partial{\mathcal{F}}_{p}(x)/\partial p$ exist.
@@
$\nabla_{p}V_{{{\mathcal{F}}_{p}}}(x_{0})$
=$\stackrel{{\scriptstyle(a)}}{{=}}\nabla_{p}\left[\sum_{u}{\mathcal{F}}_{p,u}(x_{0})\tilde{q}_{u}(x_{0},{\mathcal{F}}_{p})\right]$
=$\stackrel{{\scriptstyle(b)}}{{=}}\sum_{u}\left[\nabla_{p}{\mathcal{F}}_{p,u}(x_{0})\tilde{q}_{u}(x_{0},{\mathcal{F}}_{p})+{\mathcal{F}}_{p,u}(x_{0})\nabla_{p}\tilde{q}_{u}(x,{\mathcal{F}}_{p})\right]$
=$\stackrel{{\scriptstyle(c)}}{{=}}\sum_{u}\Bigg{[}\nabla_{p}{\mathcal{F}}_{p,u}(x_{0})\tilde{q}_{u}(x_{0},{\mathcal{F}}_{p})+\gamma{\mathcal{F}}_{p,u}(x_{0})\nabla_{p}\sum_{x^{\prime},r}\mathbb{P}_{{\mathcal{F}}_{p}}(x^{\prime},r|x,u)(r+V_{{\mathcal{F}}_{p}}(x^{\prime}))\Bigg{]}$
=$\stackrel{{\scriptstyle(d)}}{{=}}\sum_{u}\left[\nabla_{p}{\mathcal{F}}_{p,u}(x_{0})\tilde{q}_{u}(x_{0},{\mathcal{F}}_{p})+\gamma{\mathcal{F}}_{p,u}(x_{0})\sum_{x^{\prime}}\mathbb{P}_{{\mathcal{F}}_{p}}(x^{\prime}|x,u)\nabla_{p}V_{{\mathcal{F}}_{p}}(x^{\prime})\right]$
=$\stackrel{{\scriptstyle(e)}}{{=}}\sum_{u}\Bigg{[}\nabla_{p}{\mathcal{F}}_{p,u}(x_{0})\tilde{q}_{u}(x_{0},{\mathcal{F}}_{p})+\gamma{\mathcal{F}}_{p,u}(x_{0})\sum_{x^{\prime}}\mathbb{P}_{{\mathcal{F}}_{p}}(x^{\prime}|x,u)\sum_{u^{\prime}}\bigg{[}\nabla_{p}{\mathcal{F}}_{p,u^{\prime}}(x^{\prime})\tilde{q}_{u}(x^{\prime},{\mathcal{F}}_{p})$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\gamma{\mathcal{F}}_{p,u^{\prime}}(x^{\prime})\sum_{x^{\prime\prime}}\mathbb{P}_{{\mathcal{F}}_{p}}(x^{\prime\prime}|x^{\prime},u^{\prime})\nabla_{p}V_{{\mathcal{F}}_{p}}(x^{\prime\prime})\bigg{]}\Bigg{]}$
=$\stackrel{{\scriptstyle(f)}}{{=}}\sum_{x\in{{\mathbb{X}}}}\sum_{k=0}^{\infty}\gamma^{k}\left(\mathbb{P}_{{\mathcal{F}}_{p}}(x_{0},x,k)\right)\sum_{u}\nabla_{p}{\mathcal{F}}_{p,u}(x)\tilde{q}_{u}(x,{\mathcal{F}}_{p}),$
where (a) is by the definition of the value function for state $x_{0}$, (b) follows from the product rule, (c) follows by the definition of $\tilde{q}_{u}(x,{\mathcal{F}}_{p})$, (d) follows by applying the gradient w.r.t. $p$ summed over all $r$, (e) follows by repeating each of the preceding steps for $\nabla_{p}V_{{\mathcal{F}}_{p}}(x^{\prime})$, and (f) follows from repeated unrolling along the lines of (e) and upon recalling $\mathbb{P}_{{\mathcal{F}}_{p}}(x_{0},x,k)$ to be the probability of going from state $x_{0}$ to state $x$ in $k$ steps under the control policy ${\mathcal{F}}_{p}$.
@@
A.4 Control-Policy-Parameter Gradient Ascent Algorithm
Based on our new CBRL gradient ascent theorem (Theorem 2.2), we devise control-policy-parameter gradient ascent methods within the context of our general CBRL approach. One such gradient ascent method for directly learning the unknown parameter vector $p$ of the optimal control policy w.r.t. the value function $V$ comprises the iterative process according to (5) with stepsize $\eta$. Here $\nabla_{p}V_{{{\mathcal{F}}_{p}}}$ is as given by (4), where the first gradient term $\partial{\mathcal{F}}_{p,u}(x)/\partial{\mathcal{F}}_{p}$ is essentially the standard policy gradient, and the second gradient term $\partial{\mathcal{F}}_{p}(x)/\partial p$ is specific to the control-theoretic framework employed in our general CBRL approach; refer to Section 3. We note that the standard policy gradient ascent method is a special case of (5) where the parameter vector $p$ is directly replaced by the policy $\pi$; see, e.g., [21].
Our algorithmic implementation of the above control-policy-parameter gradient ascent method is summarized in Algorithm 1. This algorithm, together with the adaptation of our general CBRL approach to the LQR control-theoretic framework according to Section 3, is used to obtain the numerical results for our CBRL approach in Section 4.
@@ 4. Update the control-policy-parameter using the gradient ascent iteration (5)
Appendix B Adaptation of Control-Based Reinforcement Learning Approach
Our CBRL approach under the LQR control-theoretic framework concerns the linear dynamics $\dot{x}=Ax+Bu$ where $A$ and $B$ contain elements of the unknown parameter vector $p$. By leveraging known basic information about the LQR control problem at hand, only a relatively small number of unknown parameters need to be learned. For robotics and physics applications where state variables are derivatives of each other w.r.t. time (e.g., position, velocity, acceleration, jerk), the corresponding rows in the $A$ matrix consists of a single $1$ and the corresponding rows in $B$ comprise zeros. We exploit this basic information to consider general matrix forms for $A$ and $B$ that reduce the number of unknown parameters to be learned. As a representative illustration, the system dynamics when there are two groups of such variables have the form given by
$\underbrace{\begin{bmatrix}\frac{dx_{1}}{dt}\\ \vdots\\ \frac{d^{k_{1}}x_{1}}{dt^{k_{1}}}\\ \frac{dx_{2}}{dt}\\ \vdots\\ \frac{d^{k_{2}}x_{2}}{dt^{k_{2}}}\end{bmatrix}}_{\dot{x}}=\underbrace{\begin{bmatrix}0&1&0&\cdots&&&0\\ 0&0&1&\cdots&&\\ a_{10}&\cdots&&&&&a_{1n}\\ 0&\cdots&&0&1&\dots&0\\ 0&\cdots&&&0&1&\cdots\\ a_{20}&\cdots&&&&&a_{2n}\end{bmatrix}}_{A}x+\underbrace{\begin{bmatrix}0\\ \vdots\\ b_{0}\\ 0\\ \vdots\\ b_{1}\end{bmatrix}}_{B}u.$ (9)
Appendix C Experimental Results
In this appendix we provide additional details and results based on our numerical experiments to evaluate the performance of our CBRL approach of Section 2 adapted to the LQR control-theoretic framework according to Section 3 and using Algorithm 1. Recall that we consider the following classical control tasks from Gymnasium [28]: Cart Pole, Lunar Lander (Continuous), Mountain Car (Continuous), and Pendulum. Our CBRL approach is compared against the three state-of-the-art RL algorithms DQN [7] for discrete actions, DDPG [29] for continuous actions and PPO [23], together with a variant of PPO that solely replaces the nonlinear policy of PPO with a linear policy (since we know the optimal policy for problems such as Cart Pole is linear). These baselines are selected as the state-of-the-art RL algorithms for solving the control tasks under consideration.
Our CBRL approach depends upon the control-theoretic framework to which it is adapted, where we have chosen LQR as a representative example with the understanding that not all control tasks can be adequately solved using LQR even if the parameter vector $p$ is known. Recall, however, that our CBRL approach allows the domain of the control policies to span a subset of states in ${{\mathbb{X}}}$, thus enabling the partitioning of the state space so that properly increased richness w.r.t. finer and finer granularity can provide improved approximations and asymptotic optimality according to Theorem 2.1 of Section 2, analogous to the class of canonical piecewise-linear approximations [25]. While Cart Pole and Lunar Lander (Continuous) can be directly addressed within the context of LQR, this is not the case for Mountain Car (Continuous) and Pendulum which require a nonlinear controller. We therefore partition the state space in the case of these control tasks and consider a corresponding piecewise-LQR controller where the learned parameter vectors may differ or be shared across the partitions. Such details are provided below for Mountain Car (Continuous) and Pendulum.
C.1 Cart Pole under CBRL LQR
@@
As depicted in Figure 5, Cart Pole consists of a pole connected to a horizontally moving cart with the goal of balancing the pole by applying a force on the cart to the left or right. The state of the system is $x=[s,\dot{s},\theta,\dot{\theta}]$ in terms of the position of the cart $s$, the velocity of the cart $\dot{s}$, the angle of the pole $\theta$, and the angular velocity of the pole $\dot{\theta}$. With upright initialization of the pole on the cart, each episode comprises $200$ steps and the problem is considered “solved” upon achieving an average return of $195$.
@@
The LQR controller can be used to solve the Cart Pole problem provided that all the parameters of the system are known (e.g., mass of cart, mass and length of pole, gravity), where the angle is kept small by our CBRL approach adapted to the LQR framework. We address the problem within the context of our CBRL approach by exploiting the following general matrix form for the LQR dynamics
@@
which solely takes into account general physical relationships (e.g., the derivative of the angle of the pole is equivalent to its angular velocity) and laws (e.g., the force can only affect the acceleration) with the unknown parameters $a_{0},\ldots,a_{7},b_{0},b_{1}$ to be learned.
@@
Figure 6 presents our numerical results for the three state-of-the-art baselines (discrete actions) and our CBRL approach, each run over five independent random seeds. All parameters in (10) are initialized uniformly within $(0,1)$ and then learned using our control-policy-parameter gradient ascent iteration (5). We consider four sets of initial parameters (see Table 1) to validate the robustness of our CBRL approach. Figure 6(a) and 6(b) clearly demonstrate that our CBRL approach outperforms all other baselines w.r.t. both the number of episodes and running time, in addition to demonstrating a more stable training process. Figure 6(c) – 6(f) shows the learning behavior of CBRL parameters given different initialization.
@@
(a) Return v.s. Episode
(b) Return v.s. Time
@@
Figure 6: Learning curves of CartPole-v0 over five independent runs. The solid line shows the mean and the shaded area depicts the standard deviation. (a) and (b): Return vs. number of episodes and running time, respectively, for our CBRL approach in comparison with the Linear policy, PPO, and DQN. (c) – (f): Learning behavior of CBRL parameters, initialized by Table 1.
C.2 Lunar Lander (Continuous) under CBRL LQR
@@
As depicted in Figure 7, Lunar Lander (Continuous) is a classical spacecraft trajectory optimization problem with the goal to land softly and fuel-efficiently on a landing pad by applying thrusters to the left, to the right, and upward. The state of the system is $x=[s_{x},v_{x},s_{y},v_{y},\theta,\dot{\theta}]$ in terms of $(x,y)$ positions $s_{x}$ and $s_{y}$, two linear velocities $v_{x}$ and $v_{y}$, angle $\theta$, and angular velocity $\dot{\theta}$. Starting at the top center of the viewport with a random initial force applied to its center of mass, the lander (spacecraft) is subject to gravity, friction and turbulence while surfaces on the “moon” are randomly generated in each episode with the goal landing pad centered at $(0,0)$. The problem is considered “solved” upon achieving an average return of $200$.
@@
The LQR controller can be used to solve the Lunar Lander (Continuous) problem provided that all parameters of the system are known (e.g., mass of lander, gravity, friction, etc), where the angle is kept small by our CBRL approach adapted to the LQR framework. We address the problem within the context of our CBRL approach by exploiting the following general matrix form for the LQR dynamics
@@
which solely takes into account general physical relationships (akin to Cart Pole) and mild physical information from the system state (e.g., the acceleration is independent of the position) with the unknown parameters $a_{0},\ldots,a_{11},b_{0},b_{1},b_{2}$ to be learned.

### v2

@@ Appendix A Control-Based Reinforcement Learning Approach
Follows from Lemma 2.2 in (Hale and Cruz 1970), as given in the proof of Theorem 2 below. ∎
Intuitively, Lemma 1 shows that, for any policy families ${\mathbb{F}}_{1}\subset{\mathbb{F}}_{2}$ sufficiently close to each other, the fixed points $\tilde{q}_{1}$ and $\tilde{q}_{2}$ of the corresponding operators ${{\mathbf{T}}}_{1}$ and ${{\mathbf{T}}}_{2}$ are also close to each other. When the policy families ${\mathbb{F}}_{1}\subset{\mathbb{F}}_{2}$ are sufficiently rich and approach ${\mathbb{F}}$, then the fixed points of the corresponding operators ${{\mathbf{T}}}_{1}$ and ${{\mathbf{T}}}_{2}$ approach the unique fixed point of ${\mathbb{F}}$ satisfying (4), and therefore they approach the optimal $q$-value as promised by Bellman from Theorem 1. We formally characterize this asymptotic convergence of approximate optimality to global optimality in Theorem 2.
Theorem 2. Assume the state and action spaces are compact and ${\mathcal{F}}_{p}$ is uniformly continuous for each $p$. Consider a sequence of families of policy functions ${\mathbb{F}}_{1}\subset{\mathbb{F}}_{2}\subset\cdots\subset{\mathbb{F}}_{k-1}\subset{\mathbb{F}}_{k}$ such that $\bigcup_{i=1}^{k}{\mathbb{F}}_{i}\rightarrow{\mathbb{F}}$ as $k\rightarrow\infty$, with ${{\mathbb{S}}}$ and ${{\mathbb{S}}}_{i}$ respectively denoting the parameter sets corresponding to ${\mathbb{F}}$ and ${\mathbb{F}}_{i}$, $i\in[k]$. Then, $\sup_{x\in{{\mathbb{X}}}}\|\tilde{q}_{k}^{\prime}-\tilde{q}^{\prime}\|\rightarrow 0$ as $k\rightarrow\infty$.
Proof.
First, in the definition of the contraction mapping ${{\mathbf{T}}}$, the result of the supremum depends continuously on the set ${{\mathbb{S}}}$ and thus ${{\mathbf{T}}}$ depends continuously on ${{\mathbb{S}}}$. Next, since the state and action spaces are compact and ${\mathcal{F}}_{p}$ is uniformly continuous for each $p$, then (Hale and Cruz 1970, Lemma 2.2) shows that the unique fixed point depends continuously on the contraction mapping ${{\mathbf{T}}}$ and thus we find that $\tilde{q}^{\prime}$ depends continuously on the sets ${{\mathbb{S}}}_{i}$. This implies Lemma 1, from which Theorem 2 then follows directly. ∎
Proof of Theorem 3.
Theorem 3. Consider a family of control policy functions ${\mathbb{F}}$, its parameter set ${{\mathbb{S}}}$ with contraction operator ${{\mathbf{T}}}$ in the form of (CBRL Approach), and the value function $V_{{{\mathcal{F}}_{p}}}$ under the control policy ${\mathcal{F}}_{p}$. Assuming ${\mathcal{F}}_{p,u}(x)$ is differentiable w.r.t. ${\mathcal{F}}_{p}$ and ${\mathcal{F}}_{p}(x)$ is differentiable w.r.t. $p$ (i.e., both $\partial{\mathcal{F}}_{p,u}(x)/\partial{\mathcal{F}}_{p}$ and $\partial{\mathcal{F}}_{p}(x)/\partial p$ exist), we then have
$\nabla_{p}V_{{{\mathcal{F}}_{p}}}(x_{0})=\sum_{x\in{{\mathbb{X}}}}\>\sum_{k=0}^{\infty}\gamma^{k}\,\mathbb{P}(x,k\,|\,x_{0},{\mathcal{F}}_{p})\times\left[\sum_{u\in{\mathbb{U}}}\frac{\partial{\mathcal{F}}_{p,u}(x)}{\partial{\mathcal{F}}_{p}}\frac{\partial{\mathcal{F}}_{p}(x)}{\partial p}\tilde{q}_{u}(x,{{\mathcal{F}}}_{p})\right].$ (3)
@@
$\nabla_{p}V_{{{\mathcal{F}}_{p}}}(x_{0})$ =$\stackrel{{\scriptstyle(a)}}{{=}}\nabla_{p}\left[\sum_{u}{\mathcal{F}}_{p,u}(x_{0})\tilde{q}_{u}(x_{0},{\mathcal{F}}_{p})\right]$
=$\stackrel{{\scriptstyle(b)}}{{=}}\sum_{u}\left[\nabla_{p}{\mathcal{F}}_{p,u}(x_{0})\tilde{q}_{u}(x_{0},{\mathcal{F}}_{p})+{\mathcal{F}}_{p,u}(x_{0})\nabla_{p}\tilde{q}_{u}(x_{0},{\mathcal{F}}_{p})\right]$
=$\stackrel{{\scriptstyle(c)}}{{=}}\sum_{u}\Bigg{[}\nabla_{p}{\mathcal{F}}_{p,u}(x_{0})\tilde{q}_{u}(x_{0},{\mathcal{F}}_{p})+\gamma{\mathcal{F}}_{p,u}(x_{0})\nabla_{p}\sum_{x,r}\mathbb{P}(x,r\,|\,x_{0},{\mathcal{F}}_{p},u)(r+V_{{\mathcal{F}}_{p}}(x))\Bigg{]}$
=$\stackrel{{\scriptstyle(d)}}{{=}}\sum_{u}\left[\nabla_{p}{\mathcal{F}}_{p,u}(x_{0})\tilde{q}_{u}(x_{0},{\mathcal{F}}_{p})+\gamma{\mathcal{F}}_{p,u}(x_{0})\sum_{x}\mathbb{P}(x\,|\,x_{0},{\mathcal{F}}_{p},u)\nabla_{p}V_{{\mathcal{F}}_{p}}(x)\right]$
=$\stackrel{{\scriptstyle(e)}}{{=}}\sum_{u}\Bigg{[}\nabla_{p}{\mathcal{F}}_{p,u}(x_{0})\tilde{q}_{u}(x_{0},{\mathcal{F}}_{p})+\gamma{\mathcal{F}}_{p,u}(x_{0})\sum_{x}\mathbb{P}(x\,|\,x_{0},{\mathcal{F}}_{p},u)\sum_{u^{\prime}}\bigg{[}\nabla_{p}{\mathcal{F}}_{p,u^{\prime}}(x)\tilde{q}_{u^{\prime}}(x,{\mathcal{F}}_{p})$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\gamma{\mathcal{F}}_{p,u^{\prime}}(x)\sum_{x^{\prime}}\mathbb{P}(x^{\prime}\,|\,x,{\mathcal{F}}_{p},u^{\prime})\nabla_{p}V_{{\mathcal{F}}_{p}}(x^{\prime})\bigg{]}\Bigg{]}$
=$\stackrel{{\scriptstyle(f)}}{{=}}\sum_{x\in{{\mathbb{X}}}}\sum_{k=0}^{\infty}\gamma^{k}\,\mathbb{P}(x,k\,|\,x_{0},{\mathcal{F}}_{p})\sum_{u}\nabla_{p}{\mathcal{F}}_{p,u}(x)\tilde{q}_{u}(x,{\mathcal{F}}_{p}),$
where (a) is by the definition of the value function for state $x_{0}$, (b) follows from the product rule, (c) follows by the definition of $\tilde{q}_{u}(x,{\mathcal{F}}_{p})$, (d) follows by applying the gradient w.r.t. $p$ summed over all $r$, (e) follows by repeating each of the preceding steps for $\nabla_{p}V_{{\mathcal{F}}_{p}}(x^{\prime})$, and (f) follows from repeated unrolling along the lines of (e) and upon recalling $\mathbb{P}(x,k\,|\,x_{0},{\mathcal{F}}_{p})$ to be the probability of going from state $x_{0}$ to state $x$ in $k$ steps under the control policy ${\mathcal{F}}_{p}$.
@@
Control-Policy-Parameter Gradient Ascent Algorithm
Based on our new CBRL gradient ascent theorem (Theorem 3), we devise control-policy-parameter gradient ascent methods within the context of our general CBRL approach. One such gradient ascent method for directly learning the unknown parameter vector $p$ of the optimal control policy w.r.t. the value function $V$ comprises the iterative process according to (6) with stepsize $\eta$. Here $\nabla_{p}V_{{{\mathcal{F}}_{p}}}$ is as given by (3), where the first gradient term $\partial{\mathcal{F}}_{p,u}(x)/\partial{\mathcal{F}}_{p}$ is essentially the standard policy gradient, and the second gradient term $\partial{\mathcal{F}}_{p}(x)/\partial p$ is specific to the control-theoretic framework employed in our general CBRL approach; refer to our adaptation of the CBRL approach in the main body of the paper. We note that the standard policy gradient ascent method is a special case of (6) where the parameter vector $p$ is directly replaced by the policy $\pi$; see, e.g., (Agarwal et al. 2021).
Our algorithmic implementation of the above control-policy-parameter gradient ascent method is summarized in Algorithm 1. This algorithm, together with the adaptation of our general CBRL approach to the LQR control-theoretic framework according to the corresponding section of the main body of the paper, is used to obtain the numerical results for our CBRL approach presented in the experimental results of the main body of the paper and in the appendix below.
@@ 4. Update the control-policy-parameter using the gradient ascent iteration (6)
Appendix B Experimental Results
In this appendix we provide additional details and results based on our numerical experiments to evaluate the performance of our CBRL approach adapted to the LQR control-theoretic framework and using Algorithm 1. Recall that we consider the following classical RL tasks from Gymnasium (Towers et al. 2023): Cart Pole, Lunar Lander (Continuous), Mountain Car (Continuous), and Pendulum. Our CBRL approach is compared against the three state-of-the-art RL algorithms DQN (Mnih et al. 2013) for discrete actions, DDPG (Lillicrap et al. 2015) for continuous actions and PPO (Schulman et al. 2017), together with a variant of PPO that solely replaces the nonlinear policy of PPO with a linear policy (since we know the optimal policy for problems such as Cart Pole is linear). These baselines are selected as the state-of-the-art RL algorithms for solving the RL tasks under consideration.
Our CBRL approach depends upon the control-theoretic framework to which it is adapted, where we have chosen LQR as a representative example with the understanding that not all RL tasks can be adequately solved using LQR even if the parameter vector $p$ is known. Recall, however, that our CBRL approach allows the domain of the control policies to span a subset of states in ${{\mathbb{X}}}$, thus enabling the partitioning of the state space so that properly increased richness w.r.t. finer and finer granularity can provide improved approximations and asymptotic optimality according to Theorem 2 in the main body of the paper, analogous to the class of canonical piecewise-linear approximations (Lin and Unbehauen 1992). While Cart Pole and Lunar Lander (Continuous) can be directly addressed within the context of LQR, this is not the case for Mountain Car (Continuous) and Pendulum which require a nonlinear controller. We therefore partition the state space in the case of these RL tasks and consider a corresponding piecewise-LQR controller where the learned parameter vectors may differ or be shared across the partitions. Such details are provided below for Mountain Car (Continuous) and Pendulum.
Cart Pole under CBRL LQR
@@ Appendix B Experimental Results
As depicted in Figure 5 and described in the main body of the paper, Cart Pole consists of a pole connected to a horizontally moving cart with the goal of balancing the pole by applying a force on the cart to the left or right. The state of the system is given by $x=[s,\dot{s},\theta,\dot{\theta}]$ in terms of the position of the cart $s$, the velocity of the cart $\dot{s}$, the angle of the pole $\theta$, and the angular velocity of the pole $\dot{\theta}$.
@@
We address the Cart Pole problem within the context of our CBRL approach by exploiting the following general matrix form for the LQR dynamics
@@
which solely takes into account general physical relationships (e.g., the derivative of the angle of the pole is equivalent to its angular velocity) and laws (e.g., the force can only affect the acceleration), with the unknown parameters $a_{0},\ldots,a_{7},b_{0},b_{1}$ to be learned.
@@
Table 1 and Figure 6 present numerical results for the three state-of-the-art baselines (discrete actions) and our CBRL approach, each run over five independent random seeds. All parameters in (15) are initialized uniformly within $(0,1)$ and then learned using our control-policy-parameter gradient ascent iteration (6). We consider four sets of initial parameters (see Table 2) to validate the robustness of our CBRL approach. Figure 6(c) – 6(f) illustrates the learning behavior of CBRL parameters given different initialization.
Table 1 and Figure 6(a) and 6(b) clearly demonstrate that our CBRL approach provides far superior performance, in terms of both mean and standard deviation, over all baselines w.r.t. both the number of episodes and running time, in addition to demonstrating a more stable training process. More specifically, our CBRL approach outperforms all baselines with a significant improvement in mean return over independent random environment runs from Gymnasium (Towers et al. 2023), together with a significant reduction in the standard deviation of the return over these independent runs, thus rendering a more robust solution approach where the performance of each run is much closer to the mean than under the corresponding baseline results. In fact, all independent random environment runs under our CBRL approach yield returns equal to the mean (i.e., standard deviation of 00) for episodes $150$ and beyond. Moreover, the relative percentage difference [^1] For the relative comparison of improved mean return performance of our CBRL method over the next-best RL method, since all the performance values are positive, we use the standard relative percentage difference formula $\frac{M_{CBRL}-M_{RL}}{M_{RL}}\times 100$. Similarly, for the relative comparison of reduced standard deviation of return performance from the next-best RL method to our CBRL method, since all the performance values are positive, we typically use the standard relative percentage difference formula $\frac{S_{RL}-S_{CBRL}}{S_{CBRL}}\times 100$, with the sole exception of replacing the denominator with $(S_{RL}+S_{CBRL})/2$ based on the arithmetic mean when $S_{CBRL}$ is zero. Refer to https://en.wikipedia.org/wiki/Relative˙change in improved mean return performance of our CBRL approach over the next-best RL method Linear (based on mean performance at $500$ episodes) is $294\%$ and $6.6\%$ at $200$ and $500$ episodes, respectively; and the relative percentage difference in reduced standard deviation of return performance from the next-best RL method Linear to our CBRL approach is $200\%$ and $200\%$ at $200$ and $500$ episodes, respectively. In addition, the relative percentage difference in improved mean return performance of our CBRL approach over the DQN RL method is $927\%$ and $29.5\%$ at $200$ and $500$ episodes, respectively; and the relative percentage difference in reduced standard deviation of return performance from the DQN RL method to our CBRL approach is $200\%$ and $200\%$ at $200$ and $500$ episodes, respectively.
We note an important difference between Figure 1(a),(b) and Figure 6(a),(b), namely the shaded areas in Figure 1(a),(b) represent one form of variability across independent random seeds with the same parameter initialization, whereas the shaded areas in Figure 6(a),(b) represent the combination of two forms of variability – one across independent random seeds with the same parameter initialization and the other across different random parameter initializations.
Table 1: Mean and Standard Deviation of Return of RL Methods for CartPole-v0
Episode CBRL Linear PPO DQN
Number Mean Std. Dev. Mean Std. Dev. Mean Std. Dev. Mean Std. Dev.
50 $163.49$ $30.57$ $27.15$ $7.79$ $26.19$ $15.18$ $19.99$ $0.67$
100 $199.69$ $0.63$ $27.05$ $7.21$ $49.30$ $22.28$ $19.76$ $1.08$
150 $200.00$ $0.00$ $37.20$ $14.80$ $57.84$ $43.68$ $18.85$ $1.23$
200 $200.00$ $0.00$ $50.77$ $26.93$ $53.58$ $32.06$ $19.47$ $1.79$
250 $200.00$ $0.00$ $80.05$ $56.27$ $75.65$ $50.67$ $18.88$ $3.35$
300 $200.00$ $0.00$ $107.18$ $67.03$ $102.31$ $57.89$ $32.59$ $21.07$
350 $200.00$ $0.00$ $118.56$ $66.73$ $127.67$ $71.26$ $56.44$ $31.43$
400 $200.00$ $0.00$ $138.79$ $57.88$ $147.52$ $66.20$ $80.85$ $31.94$
450 $200.00$ $0.00$ $165.75$ $45.35$ $183.12$ $10.49$ $123.49$ $20.59$
500 $200.00$ $0.00$ $187.70$ $23.72$ $186.16$ $22.05$ $154.46$ $24.42$
@@
(a) Return vs. Episode
(b) Return vs. Time
@@
Figure 6: Learning curves of CartPole-v0 over five independent runs. The solid line shows the mean and the shaded area depicts the standard deviation. (a) and (b): Return vs. number of episodes and running time, respectively, for our CBRL approach (over the five independent runs and the four initializations in Table 2) in comparison with the Linear policy, PPO, and DQN (over the five independent runs). (c) – (f): Learning behavior of CBRL parameters, initialized by Table 2.
Lunar Lander (Continuous) under CBRL LQR
@@
As depicted in Figure 7 and described in the main body of the paper, Lunar Lander (Continuous) is a classical spacecraft trajectory optimization problem with the goal to land softly and fuel-efficiently on a landing pad by applying thrusters to the left, to the right, and upward. The state of the system is given by $x=[s_{x},v_{x},s_{y},v_{y},\theta,\dot{\theta}]$ in terms of the $(x,y)$ positions $s_{x}$ and $s_{y}$, two linear velocities $v_{x}$ and $v_{y}$, angle $\theta$, and angular velocity $\dot{\theta}$.
@@
We address the Lunar Lander (Continuous) problem within the context of our CBRL approach by exploiting the following general matrix form for the LQR dynamics
@@
which solely takes into account general physical relationships (akin to Cart Pole) and mild physical information from the system state (e.g., the acceleration is independent of the position), with the unknown parameters $a_{0},\ldots,a_{11},b_{0},b_{1},b_{2}$ to be learned.